# Improving Machine Translation with Human Feedback: An Exploration of Quality Estimation as a Reward Model

### 22 Feb 2024 v2

@@ Abstract
Insufficient modeling of human preferences within the reward model is a major obstacle for leveraging human feedback to improve translation quality. Fortunately, quality estimation (QE), which predicts the quality of a given translation without reference, has achieved impressive alignment with human evaluations in the last two years. In this work, we investigate the potential of employing the QE model as the reward model (the QE-based reward model) to predict human preferences for feedback training. We first identify the overoptimization problem during QE-based feedback training, manifested as an increase in reward while translation quality declines. We examine the problem and argue that the vulnerability of the QE model might lead to high rewards for incorrect translations, resulting in overoptimization and error propagation. To address the problem, we adopt a simple yet effective method that uses heuristic rules to detect the incorrect translations and assigns a penalty term to the QE-based rewards for the detected incorrect translations. Experimental results show that the proposed QE-based feedback training achieves consistent and significant improvements across various settings, further verified through human preference studies. Our subsequent analysis demonstrates the high data efficiency of the proposed QE-based feedback training: the proposed approach using a small amount of monolingual data can outperform systems using larger parallel corpora. Our code is available at: https://github.com/zwhe99/FeedbackMT
@@ 1 Introduction
Early efforts in neural machine translation (NMT) also attempted to integrate feedback to improve translation quality. Most works used similarity scores (such as sentence-level BLEU Papineni et al. (2002)) between the predicted translation and a reference translation to simulate feedback rather than employing feedback from humans Sokolov et al. (2016a, b); Kreutzer et al. (2017); Sokolov et al. (2017); Lawrence et al. (2017); Nguyen et al. (2017); Wu et al. (2018); Wieting et al. (2019). Few attempts to use real human feedback have been made, and these efforts either used implicit feedback in limited scenarios (e.g., e-commerce) Kreutzer et al. (2018a) or relied only on a minimal amount of human feedback data Kreutzer et al. (2018b). Therefore, the integration of real human feedback in NMT has been constrained by inadequate modeling of human preferences.
Fortunately, the field of MT has seen substantial advancements in quality estimation (QE) Rei et al. (2021, 2022b); Wan et al. (2022). A QE model offers a reference-free estimation of translation quality and has been facilitated by the growing availability of human evaluation data Specia et al. (2020, 2021); Zerva et al. (2022) and the development of pre-trained language models Devlin et al. (2019); Conneau et al. (2020). Typically, given a source sentence and its translation, a sentence-level QE model can provide a numerical score to indicate the quality of the translation. The most advanced QE models to date have achieved impressive alignment with human evaluations Freitag et al. (2022).
@@
Through manual examination of these flawed patterns, we categorize the most prevalent errors into two groups: length ratio errors and off-target errors (i.e., not the desired target language). Guided by this observation, we propose a simple yet effective solution to mitigate overoptimization. We first detect these errors and then add a negative penalty to the reward for these erroneous translations. We show that this approach significantly alleviates the overoptimization problem and results in notable improvements across various settings, which are further verified by human preference studies.
@@ 2.1 Formulation
We adopt the local ranking version of Reward rAnked FineTuning (RAFT) Dong et al. (2023) to train the model . The basic idea of RAFT is to rate the generated candidates from a prompt using the reward model and just learn from the best one of them. It has been proved to be more stable and efficient than Proximal Policy Optimization (PPO) Schulman et al. (2017). Algorithm 1 shows the details of RAFT.
@@ 2.2 Addressing Overoptimization Problem
We further found that the reward model may assign high scores to erroneous translations in some cases (as shown in Table 1). These errors encompass common error patterns in MT, such as length-ratio, off-target errors, and hallucinations111For sake of simplicity, we categorize repeating errors as length-ratio errors, and copying-source errors are considered as one kind of off-target errors..
@@
While these errors may not be severe initially, they can rapidly propagate to subsequent training stages (see Figure 2) once they are accorded high reward scores, which can lead to disruption of the entire training process.
@@ 3.1 Experimental Setup
We adopt the pipeline of reinforcement learning with human feedback (RLHF) Ouyang et al. (2022). Starting with the pretrained base model, we carry out the following steps: (1) Supervised fine-tuning (SFT), where we utilize parallel data to fine-tune the pretrained base model, thereby obtaining an initial MT model. (2) Feedback training, where we train the model using RAFT/RAFT+ to maximize the reward from a QE-based reward model. Note that this stage only uses monolingual data.
@@
We consider the settings for both high-resource and low-resource language pairs. In the high-resource setting, we focus on EnglishChinese and EnglishGerman. In the low-resource setting, our focus is on EnglishUkrainian and UkrainianCzech. It is important to note that both settings are multilingual, where all four translation directions are trained concurrently within one model, and we ensure the balance of all directions to avoid introducing other factors. That is, every direction has an equal number of training samples. We adopt WMT22 Kocmi et al. (2022) as the test sets for all language pairs. For development sets, high resource setting uses WMT21, and low resource setting uses Flores200 Goyal et al. (2022). Table 2 lists the detail of the data.
@@ 3.2 Results
Feedback training brings significant improvements in general. Regardless of whether the resource setting is high or low and irrespective of the variation in base models, RAFT+ tends to deliver notable enhancements, particularly pronounced in the low resource setting. This suggests that current QE models are already equipped with the capability to function as reward models after addressing the overoptimization problem. In contrast, while achieving good gains in most settings, RAFT can suffer from severe overoptimization that might lead to failed training.
@@
Remarkably, even COMET, a reference-based metric, can be overoptimized. In RAFT training of Nllb-200-1.3B under low-resource setting (bottom of Table 2(b)), there is a marginal increase in COMET scores, yet a significant drop in BLEURT, which is unusual. By inspecting the outputs, we found that severe off-target errors significantly hamper the model’s performance. One plausible reason might be the similarities between the COMET and COMET-QE models, both of which may be susceptible to these off-target errors. This finding underscores the necessity of treating automatic metrics with caution and the importance of employing multiple metrics simultaneously for a more comprehensive evaluation. It also stresses the significance of recognizing a metric model’s vulnerabilities during its development, including adjustments for prevalent translation inaccuracies such as length-ratio, off-target errors, hallucinations, and so forth.
@@ 4.2 Data Efficiency of Feedback Training
Data efficiency refers to the capacity to achieve good performance with less training data. For low-resource language pairs, data efficiency is important since it is costly and labor-intensive to annotate high-quality parallel data. Feedback training provides an alternative path where humans only need to evaluate rather than generate translations.
@@ 4.4 Effects of Hyperparameters
The sampling temperature and the number of candidates are two crucial hyperparameters for RAFT+. Both influence the diversity of the candidates. Intuitively, having a broader range of candidates may provide richer feedback signals, potentially leading to better final performance. However, as illustrated in Figure 7, (1) continuously increasing the temperature leads to performance degradation; (2) the boost from continuously increasing the number of candidates approaches saturation. This means that the increase in candidate diversity has reached its upper limit.
@@ 5.2 Aligning LLMs with Human
Pretrained on raw text from the internet, LLMs might generate toxic, inaccurate, and unhelpful content Fernandes et al. (2023). To mitigate these issues, researchers have employed human feedback to better align the behavior of LLMs with human preferences, thereby enhancing their helpfulness and reducing potential harm Ouyang et al. (2022); Stiennon et al. (2020); Bai et al. (2022). The basic idea involves learning a reward function that capturing human preferences and optimizing the LLM using proximal policy optimization (PPO) Schulman et al. (2017). Lighter weight training strategies, such as RAFT Dong et al. (2023) and RRHF Yuan et al. (2023), use training data ranking based on rewards to align the LLM. LIMA Zhou et al. (2023) underscores the importance of high-quality supervised data for effective alignment. For a comprehensive understanding of alignment, we recommend readers refer to survey papers: Fernandes et al. (2023) and Wang et al. (2023).
@@ 6 Conclusion
This work only considers a simple yet stable training algorithm, RAFT, in favor of its ease of use and reduced computational requirements. We also implemented the Minimum Risk Training (MRT) algorithm Shen et al. (2016) in Appendix B, but found it difficult to train stably. While other more commonly used reinforcement learning (RL) algorithms such as Proximal Policy Optimization (PPO) Schulman et al. (2017) exist, their training stability can be influenced by a variety of factors and often demands substantial GPU memory Dong et al. (2023); Zheng et al. (2023). Hyperparameters, stability, and computational constraints led us to consciously limit our exploration to RAFT, thereby not investigating other potentially effective but more complex training algorithms. This choice simplifies the training process but might limit the general applicability of the approach in different contexts. Future work may explore these alternative algorithms, acknowledging the trade-offs between complexity, stability, and performance.
@@
This work does not consider the imbalances present in multilingual NMT. In real-world scenarios, the quantity of available corpus varies significantly across different languages, which might lead to biased or suboptimal results. However, considering unbalanced scenarios introduces many additional factors, such as different ways of sampling training data. Therefore, we constrain our focus to the balanced situation in this work, recognizing the need for future research to address this complexity and explore methods that can better handle the imbalances inherent in multilingual contexts.
@@
Although we kept the amount of data the same for the different language pairs, we still observed training imbalance problems due to “uneven” distribution of rewards. Specifically, in Table 2(a), neither RAFT nor RAFT+ consistently improves NLLB in EnDe and EnZh (both are from-English direction) at the high resource setting. However, on the other hand, the performance of the other two to-English directions (DeEn and ZhEn) has been significantly improved by notable margins (BLEURT | DeEn: 53.571.3, ZhEn: 48.462.9). This means that during training, the model allocates more “capacity” to the to-English directions than to the from-English directions, i.e., they are not balanced. The intuitive reason is that the model “finds” optimizing the to-English directions offers a quick boost in reward, thereby “ignoring” the from-English directions which have less room for improvement. This involves how the reward can be reasonably distributed among different language pairs, which is less discussed in this work, but a meaningful and underexplored future direction.
@@ Appendix A Error Detection and Penalty
We penalize for len-ratio errors and off-target errors. We consider translations whose length ratios do not belong to as len-ratio errors. We computed the distribution of length ratios of the SFT data for each language pair (using the tokenizer of the corresponding model), taking so that it covers the top 50% of the most frequent length ratios. We adopt language detector from Stahl (2023) to detect off-target errors.
@@ Appendix B MRT Implementation
MRT Shen et al. (2016) is designed to directly optimize the model with respect to evaluation metrics. The risk of MRT can be formulate as:
@@
and is a hyperparameter. Algorithm 2 shows the details of MRT. Similar to RAFT+, when adding a penalty term to reward, we call it MRT+. We set the penalty term and 222We follow the default setting in Sennrich et al. (2017)., and leave other hyperparameters the same as RAFT.
@@ Appendix D UniTE-MUP as the Reward Model
| Method | DeEn | EnDe | ZhEn | EnZh | Average | |||||
| Method | EnUk | UkEn | UkCs | CsUk | Average | |||||

### 18 Mar 2024 v3

@@ Abstract
Insufficient modeling of human preferences within the reward model is a major obstacle for leveraging human feedback to improve translation quality. Fortunately, quality estimation (QE), which predicts the quality of a given translation without reference, has achieved impressive alignment with human evaluations in the last two years. In this work, we investigate the potential of employing the QE model as the reward model to predict human preferences for feedback training. We first identify the overoptimization problem during QE-based feedback training, manifested as an increase in reward while translation quality declines. We examine the problem and argue that the vulnerability of the QE model might lead to high rewards for incorrect translations, resulting in overoptimization and error propagation. To address the problem, we adopt a simple yet effective method that uses heuristic rules to detect the incorrect translations and assigns a penalty term to the reward scores of them. Experimental results show that the proposed QE-based feedback training achieves consistent and significant improvements across various settings, further verified through human preference studies. Our subsequent analysis demonstrates the high data efficiency of the proposed QE-based feedback training: it outperforms systems using larger parallel corpora by a small amount of monolingual data. Our code is available at: https://github.com/zwhe99/FeedbackMT
@@ 1 Introduction
Early efforts in neural machine translation (NMT) also attempted to integrate feedback to improve translation quality. Most works used similarity scores (such as sentence-level BLEU; Papineni et al. 2002) between the predicted translation and a reference translation to simulate feedback rather than employing feedback from humans Sokolov et al. (2016a, b); Kreutzer et al. (2017); Sokolov et al. (2017); Lawrence et al. (2017); Nguyen et al. (2017); Wu et al. (2018); Wieting et al. (2019). Few attempts to use real human feedback have been made, and these efforts either used implicit feedback in limited scenarios (e.g., e-commerce; Kreutzer et al. 2018a) or relied only on a minimal amount of human feedback data Kreutzer et al. (2018b). Therefore, the integration of real human feedback in NMT has been constrained by inadequate modeling of human preferences.
Fortunately, the field of MT has seen substantial advancements in quality estimation (QE; Rei et al. 2021, 2022b; Wan et al. 2022). A QE model offers a reference-free estimation of translation quality and has been facilitated by the growing availability of human evaluation data Specia et al. (2020, 2021); Zerva et al. (2022) and the development of pre-trained language models Devlin et al. (2019); Conneau et al. (2020). Typically, given a source sentence and its translation, a sentence-level QE model can provide a numerical score to indicate the quality of the translation. The most advanced QE models to date have achieved impressive alignment with human evaluations Freitag et al. (2022) and have been used to guide the decoding process Fernandes et al. (2022); He et al. (2024).
@@
Through manual examination of these flawed patterns, we categorize the most prevalent errors into two groups: length ratio errors and off-target errors (i.e., not the desired target language). Guided by this observation, we propose a simple yet effective solution to mitigate overoptimization. We first detect these errors and then add a negative penalty to the reward for these erroneous translations. We show that this approach significantly alleviates the overoptimization problem and results in notable improvements across various settings, which are further verified by human preference studies111Automatic metrics and human judgement might be not consistent under feedback training as discussed in Limitations.
@@ 2.1 Formulation
We adopt the local ranking version of Reward rAnked FineTuning (RAFT; Dong et al. 2023) to train the model . The basic idea of RAFT is to rate the generated candidates from a prompt using the reward model and just learn from the best one of them. It has been proved to be more stable and efficient than Proximal Policy Optimization (PPO; Schulman et al. 2017). Algorithm 1 shows the details of RAFT.
@@ 2.2 Addressing Overoptimization Problem
We further found that the reward model may assign high scores to erroneous translations in some cases (as shown in Table 1). These errors encompass common error patterns in MT, such as length-ratio, off-target errors, and hallucinations222For simplicity, we categorize repeating errors as length-ratio errors, and consider copying-source errors under the category of off-target errors..
@@
While these errors may not be severe initially, they can rapidly propagate to subsequent training stages (see Figure 2) once they are accorded high reward scores, which can lead to disruption of the entire training process. Yan et al. (2023) also observed a similar phenomenon when optimizing BLEURT.
@@ 3.1 Experimental Setup
We adopt the pipeline of reinforcement learning with human feedback (RLHF; Ouyang et al. 2022). Starting with the pretrained base model, we carry out the following steps: (1) Supervised fine-tuning (SFT), where we utilize parallel data to fine-tune the pretrained base model, thereby obtaining an initial MT model. (2) Feedback training, where we train the model using RAFT/RAFT+ to maximize the reward from a QE model. Note that this stage only uses monolingual data.
@@
We consider the settings for both high- and low-resource language pairs. In the high-resource setting, we focus on EnglishChinese and EnglishGerman. In the low-resource setting, our focus is on EnglishUkrainian and UkrainianCzech. It is important to note that both settings are multilingual, where all four translation directions are trained concurrently within one model, and we ensure the balance of all directions to avoid introducing other factors. That is, every direction has an equal number of training samples. We adopt WMT22 Kocmi et al. (2022); He et al. (2022a) as the test sets for all language pairs, which is latest enough to avoid data contamination Zhu et al. (2023b, a). For development sets, high resource setting uses WMT21, and low resource setting uses Flores200 Goyal et al. (2022). Table 6 lists the detail of the data we used.
@@ 3.2 Results
Feedback training brings significant improvements in general. Regardless of whether the resource setting is high or low and irrespective of the variation in base models, RAFT+ tends to deliver notable enhancements (though still under-performs the WMT22 Best systems), particularly pronounced in the low resource setting. This suggests that current QE models are already equipped with the capability to function as reward models after addressing the overoptimization problem. In contrast, while achieving good gains in most settings, RAFT can suffer from severe overoptimization that might lead to failed training.
@@
Remarkably, even COMET, a reference-based metric, can be overoptimized. In RAFT training of Nllb-200-1.3B under low-resource setting (bottom of Table 2(b)), there is a marginal increase in COMET scores, yet a significant drop in BLEURT, which is unusual. By inspecting the outputs, we found that severe off-target errors significantly hamper the model’s performance. One plausible reason might be the similarities between the COMET and COMET-QE models, both of which may be susceptible to these off-target errors. This finding underscores the necessity of treating automatic metrics with caution and the importance of employing multiple metrics simultaneously for a more comprehensive evaluation. Especially for feedback training using a reward model, the increasement of automatic metrics might not indicate true translation quality improvement. We therefore conducted a human evaluation in § 4.1 and have a discussion in Limitations. It also stresses the significance of recognizing a metric’s vulnerabilities during its development, including adjustments for prevalent translation inaccuracies such as length-ratio, off-target errors, hallucinations.
@@ 4.2 Data Efficiency of Feedback Training
Data efficiency refers to the capacity to achieve good performance with less training data. For low-resource language pairs, data efficiency is important since it is costly to annotate high-quality parallel data He et al. (2022b). Feedback training provides an alternative path where humans only need to evaluate rather than generate translations.
@@ 4.4 Effects of Hyperparameters
The sampling temperature and the number of candidates are two crucial hyperparameters for RAFT+. Both influence the diversity of the candidates. Intuitively, having more diverse candidates may provide richer feedback signals, potentially leading to better final performance. However, as illustrated in Figure 7, (1) continuously increasing the temperature leads to performance degradation; (2) the boost from continuously increasing the number of candidates approaches saturation. This means that the increase in candidate diversity has reached its upper limit.
@@ 5.1 Feedback Training in MT
Kreutzer et al. (2018a) utilizes implicit human feedback in a constrained e-commerce scenario, which cannot be applied to general MT. Kreutzer et al. (2018b) investigates how to improve translation using human evaluation data but is limited by the small data size (1K). Concurrently with this work, Finkelstein and Freitag (2023); Ramos et al. (2023) and Gulcehre et al. (2023) also consider using QE models to provide reward signals in the training of MT models. Xu et al. (2024) points out that the references of training data might be worse than model outputs in quality, which aligns with our findings in § 4.2.
@@ 5.2 Aligning LLMs with Human
Pretrained on raw text from the internet, LLMs might generate toxic, inaccurate, and unhelpful content Fernandes et al. (2023). To mitigate these issues, researchers have employed human feedback to better align the behavior of LLMs with human preferences, thereby enhancing their helpfulness and reducing potential harm Ouyang et al. (2022); Stiennon et al. (2020); Bai et al. (2022). The basic idea involves learning a reward function that captures human preferences and optimizing the LLM using proximal policy optimization (PPO; Schulman et al. 2017) to maximize the reward. Lighter weight training strategies, such as RAFT Dong et al. (2023), RRHF Yuan et al. (2023) and DPO Rafailov et al. (2024), use training data ranking based on rewards or human annotations to align the LLM. LIMA Zhou et al. (2023) underscores the importance of high-quality supervised data for effective alignment. For a comprehensive understanding of alignment, we recommend readers refer to survey papers: Fernandes et al. (2023) and Wang et al. (2023).
@@ 6 Conclusion
Acknowledgements
Zhiwei and Rui are with MT-Lab, Department of Computer Science and Engineering, School of Electronic Information and Electrical Engineering, and also with the MoE Key Lab of Artificial Intelligence, AI Institute, Shanghai Jiao Tong University, Shanghai 200204, China. Rui and Zhiwei are supported by the Tencent Open Fund (RBFR2023012), the National Natural Science Foundation of China (62176153), and the Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102).
@@ Acknowledgements
An increase in automatic metrics do not necessarily indicate a true improvement in translation quality.
When we conducting feedback training, we need a QE model to provide reward score and a metric model to evaluate the final performance. Though they play different roles in the whole pipeline, they share the same function of assessing translation quality. Unfortunately, current open-sourcing QE and metric models that show strong correlation with human mostly rely on human evaluation data in WMT, which indicating the overlap between their training data. Therefore, overoptimizing the QE model might lead to unconsciously overoptimizing the metric model, as we observed in § 3.2. This is the reason why we conducted a human preference study to double check the improvement. However, human annotation is always costly and not scalable, and we only conduct it on EnZh.
@@
This work only considers a simple yet stable training algorithm, RAFT, in favor of its ease of use and reduced computational requirements. We also implemented the Minimum Risk Training (MRT) algorithm Shen et al. (2016) in Appendix B, but found it difficult to train stably. While other more commonly used reinforcement learning (RL) algorithms such as PPO exist, their training stability can be influenced by a variety of factors and often demands substantial GPU memory Dong et al. (2023); Zheng et al. (2023). Hyperparameters, stability, and computational constraints led us to consciously limit our exploration to RAFT, thereby not investigating other potentially effective but more complex training algorithms. This choice simplifies the training process but might limit the general applicability of the approach in different contexts. Future work may explore these alternative algorithms, acknowledging the trade-offs between complexity, stability, and performance.
@@
This work does not consider the imbalances present in multilingual NMT. In real-world scenarios, the quantity of available corpora varies significantly across different languages, which might lead to biased or suboptimal results. However, considering unbalanced scenarios introduces many additional factors, such as different ways of sampling training data. Therefore, we constrain our focus to the balanced situation in this work, recognizing the need for future research to address this complexity and explore methods that can better handle the imbalances inherent in multilingual contexts.
@@
Although we kept the amount of data the same for different language pairs, we still observed training imbalance problems due to “uneven” distribution of rewards. Specifically, in Table 2(a), neither RAFT nor RAFT+ consistently improves NLLB in EnDe and EnZh (both are from-English direction) at the high resource setting. On the other hand, the performance of the other two to-English directions (DeEn and ZhEn) has been significantly improved by notable margins (BLEURT | DeEn: 53.571.3, ZhEn: 48.462.9). This means that during training, the model allocates more “capacity” to the to-English directions than to the from-English directions, i.e., they are not balanced. The intuitive reason is that the model “finds” optimizing the to-English directions offers a quick boost in reward, thereby “ignoring” the from-English directions which have less room for improvement. This involves how the reward can be reasonably distributed among different language pairs, which is less discussed in this work, but a meaningful and underexplored future direction.
@@ Appendix A Error Detection and Penalty
We penalize for len-ratio errors and off-target errors. We consider translations whose length ratios do not belong to as len-ratio errors. We computed the distribution of length ratios of the SFT training data for each language pair (using the tokenizer of the corresponding model), taking so that it covers the top 50% of the most frequent length ratios. We adopt language detector from Stahl (2023) to detect off-target errors.
@@ Appendix B MRT Implementation
MRT Shen et al. (2016) is designed to directly optimize the model with respect to evaluation metrics. However, we failed to get meaningful improvements when using MRT.
The risk of MRT can be formulate as:
@@
and is a hyperparameter. Algorithm 2 shows the details of MRT. Similar to RAFT+, when adding a penalty term to reward, we call it MRT+. We set the penalty term and 777We follow the default setting in Sennrich et al. (2017)., and leave other hyperparameters the same as RAFT.